1a) company T stock has just paid its most recent dividend of $2.00. the year end dividends are expected to decline at 5% per year forever. with a required return of 14%, what is the expected price in 2 years?

b) what is the expected price today ?
Answer #1

D1= 2*95% = 1.90

D2 = 1.90*95% = 1.81

D3 = 1.81*95% = 1.71

price at time 2 = D3/(k-g) = 1.71/(14%-(-5%))

= 9.03

Price today = D1/(k-g)

=1.9/(14%+5%)

= $10